You get a box of assorted chocolates for Christmas. There are the following
types; 8 dark chocolate-coated caramels, 7 milks chocolate-coated caramels,7
dark chocolate-coated nuts, 6 milk chocolate-coated vanilla creams, 8 dark
chocolate-coated cherries, 7 milk chocolate-coated strawberry creams. Determine
the following probabilitie;
of getting 2 dark chocolate-coated cherries in a row and then 5
caramels in a row.''.

Answer by venugopalramana(3286)

There are the following
types;
8 dark chocolate-coated caramels......8DCCC,
7 milks chocolate-coated caramels.........7MCCC
7 dark chocolate-coated nuts......7DCCN,
6 milk chocolate-coated vanilla creams.......6MCCVC,
8 dark chocolate-coated cherries.....8DCCCH,
7 milk chocolate-coated strawberry creams.......7MCCSC.
TOTAL =43...
Determine the following probabilitY;
of getting 2 dark chocolate-coated cherries....2DCCCH in a row
and then 5 caramels ...IT IS ASSUMED HERE THAT THEY ARE DARK CHACOLATE COATED CARAMELS...5DCCC in a row.''. (ANY WAY THE METHOD WILL GUIDE YOU TO DO IF THEY ARE MILK CHACOLATE COATED CARAMELS)
TOTAL =43..WE HAVE TO DRAW THE CHACOLATES ONLY IN THE GIVEN ORDER.HENCE
I DRAW ..........TOTAL =43...DCCCH=8...............PROBABILITY =8/43
II DRAW..........TOTAL =42...DCCCH=7...............PROBABILITY =7/42
III DRAW.........TOTAL =41...DCCC =8...............PROBABILITY =8/41
1V DRAW..........TOTAL =40...DCCC =7...............PROBABILITY =7/40
V DRAW...........TOTAL =39...DCCC =6...............PROBABILITY =6/39
VI DRAW..........TOTAL =38...DCCC =5...............PROBABILITY =5/38
VII DRAW.........TOTAL =37...DCCC =4...............PROBABILITY =4/37
SO THE COMPOUND PROBABILITY IS THE PRODUCT OF ALL THE ABOVE PROBABILITIES...
(8*7*8*7*6*5*4)/(43*42*41*40*39*38*37)